# IL-SOAR : Imitation Learning with Soft Optimistic Actor cRitic

### v2

@@ Abstract
When instantiated in the tabular setting, we get a provable algorithm with guarantees that matches the best known results in $\epsilon$.
Practically, the SOAR template is shown to boost consistently the performance of imitation learning algorithms based on Soft Actor Critic such as $f$-IRL, ML-IRL and CSIL in several MuJoCo environments. Overall, thanks to SOAR, the required number of episodes to achieve the same performance is reduced by half.
@@ 3 The Algorithm
Notice that only one critic is used in the implementation of SAC ($L=1$) that serves as base RL algorithm for several commonly used IL algorithms ( GAIL [28], AIRL [19], IQ-Learn [24], PPIL [98], RKL [52] and ML-IRL [111]). As proven in Corollary 4.11, $L=1$ is not enough to ensure optimism, not even in the tabular case. In our experiments, we show that a value of $L$ larger than $1$ is beneficial in all the MuJoCo environments we tested on.
@@ 4 Theoretical analysis
Consider the MDP $M=(\mathcal{S},\mathcal{A},\gamma,P,c)$ and two policies $\pi,\pi^{\prime}:\mathcal{S}\rightarrow\Delta_{\mathcal{A}}$. Then consider for any $\widehat{Q}\in\mathbb{R}^{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|}$ and $\widehat{V}^{\pi}(s)=\left\langle{\pi(\cdot|s)},{\widehat{Q}(s,\cdot)}\right\rangle$ and $Q^{\pi^{\prime}},V^{\pi^{\prime}}$ be respectively the state-action and state value function of the policy $\pi$ in MDP $M$. Then, it holds that $(1-\gamma)\left\langle{\boldsymbol{\nu}_{0}},{\widehat{V}^{\pi}-V^{\pi^{\prime}}}\right\rangle$ equals
@@
Consider an MDP where $\max_{s,a\in\mathcal{S}\times\mathcal{A}}\mathrm{supp}(P(\cdot|s,a))=2$, then for $L=36\log\left({\frac{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|K}{\delta}}\right)$ it holds that with probability at least $1-\delta$
@@
To fix this issue, we consider multiple critics and we used as an optimistic estimate the mean minus the standard deviation of the ensemble as explained in Algorithm 5. In addition, the standard deviation needs to be truncated at a threshold, as was done in the tabular analysis, to avoid the value function estimators growing out of the attainable range. For any state $s$, the attainable range for the state value function is $\left[{0,(1-\gamma)^{-1}}\right]$.
@@ Appendix A Related Works
Our theoretical guarantees are in the same setting of previous works like [82] and [106] with the difference that their work focuses on the easier finite horizon setting. Additionally, their exploration techniques only applies to the tabular setting. Indeed, in the MuJoCo experiments in [82], the authors do not attempt to implement the exploration mechanism required for their theoretical guarantees. In a similar way, also [106] can not be implemented beyond the tabular setting because it relies on a reward free procedure requiring bonuses proportional to the number of visits to each state action pair. [76] suggest an algorithm which does not require exploration but it can not improve upon behavioural cloning in terms of expert trajectories. [71, 73, 18] analyze instead offline imitation learning (behavioural cloning) where no additional interaction with the environment is allowed. This setting is more general but it comes at the cost of additional assumptions such as policy realizability or worst depedence on the horizon on the required number of trajectories. [18] presents an analysis for general policy classes but they require a maximum likelihood oracle which can not be implemented exactly when using neural network function approximation.
@@
As mentioned only few works investigated exploration techniques in Deep IL. Apart from the previously mentioned works, [110] adopts a model based approach and used exploration bonuses based on prediction error of the next observed state (a.k.a. curiosity driven exploration [65, 6]).
@@ B.1 Upper bounding the policy regret
where the last inequality holds for $N^{k}_{\ell}\geq 2$. Otherwise, for $P(s^{\prime}|s,a)\leq 1-\frac{1}{N^{k}_{\ell}(s,a)}$ we can apply [8, Lemma 2] (adapted from [103, Corollary 1] ) to obtain
@@ B.2 Policy Regret Decomposition
$\sum^{K}_{k=1}\sum_{s,a}\left[{d^{\pi^{k}}(s,a)-d^{\pi^{\star}}(s,a)}\right]\cdot\left[{\delta^{k}(s,a)}\right]$ $\leq\gamma\sum^{K}_{k=1}\mathbb{E}_{s,a\sim d^{\pi^{k}}}\left[{\sqrt{\frac{\left|{\mathcal{S}}\right|}{(N^{k}(s,a)/L+1)(1-\gamma)^{2}}\log\left({\frac{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|LK}{(1-\gamma)\delta}}\right)}}\right]$
@@
$\leq\sqrt{K\sum^{K}_{k=1}\mathbb{E}_{s,a\sim d^{\pi^{k}}}\left[{\frac{\left|{\mathcal{S}}\right|}{(N^{k}(s,a)/L+1)(1-\gamma)^{2}}\log\left({\frac{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|LK}{(1-\gamma)\delta}}\right)}\right]}$
@@
$\sum^{K}_{k=1}\sum_{s,a}\left[{d^{\pi^{k}}(s,a)-d^{\pi^{\star}}(s,a)}\right]\cdot\left[{\delta^{k}(s,a)}\right]$ $\leq\frac{1}{1-\gamma}\sqrt{K\left|{\mathcal{S}}\right|\log\left({\frac{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|LK}{(1-\gamma)\delta}}\right)(2L\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|\log\left({KL_{\max}}\right)+4\log\left({2K/\delta}\right))}$
@@
$\leq\left({\sqrt{\frac{\left|{\mathcal{S}}\right|}{(N^{k}(s,a)/L+1)(1-\gamma)^{2}}\log\left({\frac{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|LK}{(1-\gamma)\delta}}\right)}+\frac{2}{K}+\frac{2}{(N^{k}(s,a)/L+1)(1-\gamma)}}\right)^{2}$
$\leq\frac{3\left|{\mathcal{S}}\right|}{(N^{k}(s,a)/L+1)(1-\gamma)^{2}}\log\left({\frac{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|LK}{(1-\gamma)\delta}}\right)+\frac{12}{K^{2}}+\frac{12}{(N^{k}(s,a)/L+1)^{2}(1-\gamma)^{2}}$
@@
$PV^{k}(s,a)-\widehat{P}^{k}_{\ell}V^{k}(s,a)$ $\leq\sqrt{\frac{\left|{\mathcal{S}}\right|}{(N^{k}(s,a)/L+1)(1-\gamma)^{2}}\log\left({\frac{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|LK}{(1-\gamma)\delta}}\right)}+\frac{2}{K}$
$\phantom{=}+\frac{2}{(N^{k}(s,a)/L+1)(1-\gamma)}~{}~{}~{}~{}~{}\forall\ell\in[L]$
In particular, the above statement implies that
$PV^{k}(s,a)-\min_{\ell\in[L]}\widehat{P}^{k}_{\ell}V^{k}(s,a)$ $\leq\sqrt{\frac{\left|{\mathcal{S}}\right|}{(N^{k}(s,a)/L+1)(1-\gamma)^{2}}\log\left({\frac{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|LK}{(1-\gamma)\delta}}\right)}+\frac{2}{K}$
@@
Then, by Azuma-Hoeffding inequality it holds that with probability $1-\delta$
$\sum_{\bar{s},\bar{a},s^{\prime}\in\mathcal{R}^{k}_{\ell}}\left({P\tilde{V}(s,a)-\tilde{V}(s^{\prime})}\right)\mathds{1}_{\left\{{s,a=\bar{s},\bar{a}}\right\}}\leq\sqrt{\frac{N^{k}_{\ell}(s,a)\log(1/\delta)}{(1-\gamma)^{2}2}}.$
Therefore, we can conclude that with probability at least $1-\delta$
$P\tilde{V}(s,a)-\frac{1}{N^{k}_{\ell}(s,a)}\sum_{\bar{s},\bar{a},s^{\prime}\in\mathcal{R}^{k}_{\ell}}\tilde{V}(s^{\prime})\mathds{1}_{\left\{{s,a=\bar{s},\bar{a}}\right\}}\leq\sqrt{\frac{\log(1/\delta)}{N^{k}_{\ell}(s,a)(1-\gamma)^{2}2}}.$
Now, by a union bound over $\mathcal{C}_{\epsilon_{\mathrm{cov}}}(\mathcal{V})$ and $[L]$ and $\mathcal{S}\times\mathcal{A}$ and denoting
@@
$\mathbb{P}\left[{P\tilde{V}(s,a)-\bar{P}^{k}_{\ell}\tilde{V}(s,a)\leq\sqrt{\frac{\log(\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|\left|{\mathcal{C}_{\epsilon_{\mathrm{cov}}}(\mathcal{V})}\right|L/\delta)}{N^{k}_{\ell}(s,a)(1-\gamma)^{2}}}~{}~{}\forall s,a,\ell\in\mathcal{S}\times\mathcal{A}\times[L],~{}~{}\tilde{V}\in\mathcal{V}}\right]\geq 1-\delta$
@@
$\leq\sqrt{\frac{\log(\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|\left|{\mathcal{C}_{\epsilon_{\mathrm{cov}}}(\mathcal{V})}\right|L/\delta)}{N^{k}_{\ell}(s,a)(1-\gamma)^{2}}}+2\epsilon_{\mathrm{cov}}$
$\leq\sqrt{\frac{\left|{\mathcal{S}}\right|}{N^{k}_{\ell}(s,a)(1-\gamma)^{2}}\log\left({\frac{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|L}{(1-\gamma)\epsilon_{\mathrm{cov}}\delta}}\right)}+2\epsilon_{\mathrm{cov}}$
@@
$PV^{k}(s,a)-\bar{P}^{k}_{\ell}V^{k}(s,a)\leq\sqrt{\frac{\left|{\mathcal{S}}\right|}{N^{k}_{\ell}(s,a)(1-\gamma)^{2}}\log\left({\frac{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|LK}{(1-\gamma)\delta}}\right)}+\frac{2}{K}$
@@
$\geq\frac{N^{k}_{\ell}(s,a)}{N^{k}_{\ell}(s,a)+2}\left[{PV^{k}(s,a)-\sqrt{\frac{\left|{\mathcal{S}}\right|}{N^{k}_{\ell}(s,a)(1-\gamma)^{2}}\log\left({\frac{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|LK}{(1-\gamma)\delta}}\right)}-\frac{2}{K}}\right]$
$=PV^{k}(s,a)-\frac{2}{N^{k}_{\ell}(s,a)+2}PV^{k}(s,a)-\sqrt{\frac{\left|{\mathcal{S}}\right|}{(N^{k}_{\ell}(s,a)+2)(1-\gamma)^{2}}\log\left({\frac{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|LK}{(1-\gamma)\delta}}\right)}-\frac{2}{K}$
$\geq PV^{k}(s,a)-\frac{2}{(N^{k}_{\ell}(s,a)+2)(1-\gamma)}-\sqrt{\frac{\left|{\mathcal{S}}\right|}{(N^{k}_{\ell}(s,a)+2)(1-\gamma)^{2}}\log\left({\frac{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|LK}{(1-\gamma)\delta}}\right)}-\frac{2}{K}$
Finally, rearranging and using that $N^{k}_{\ell}(s,a)=\left\lfloor\frac{N^{k}(s,a)}{L}\right\rfloor\geq\frac{N^{k}(s,a)}{L}-1,$ we obtain that with probability $1-\delta$, it holds that for all $\ell\in[L]$
$PV^{k}(s,a)-\widehat{P}^{k}_{\ell}V^{k}(s,a)$ $\leq\sqrt{\frac{\left|{\mathcal{S}}\right|}{(N^{k}(s,a)/L+1)(1-\gamma)^{2}}\log\left({\frac{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|LK}{(1-\gamma)\delta}}\right)}+\frac{2}{K}$

### v3

@@ Abstract
When instantiated in the tabular setting, we get a provable algorithm with guarantees that matches the best known results in the desired accuracy parameter $\epsilon$.
Practically, the SOAR template can boost the performance of any imitation learning algorithm based on Soft Actror Critic (SAC). As an example, we show that SOAR can boost consistently the performance of the following SAC-based imitation learning algorithms: $f$-IRL, ML-IRL and CSIL. Overall, thanks to SOAR, the required number of episodes to achieve the same performance is reduced by half. [^1] Project code available at https://github.com/stefanoviel/SOAR-IL/tree/master
@@ 3 The Algorithm
Notice that only one pair of critics is used in the implementation of SAC ($L=1$) that serves as base RL algorithm for several commonly used IL algorithms ( GAIL [30], AIRL [21], IQ-Learn [26], PPIL [102], RKL [55] and ML-IRL [115]). As proven in Corollary 4.11, $L=1$ is not enough to ensure optimism, not even in the tabular case. In our experiments, we show that a value of $L$ larger than $1$ is beneficial in all the MuJoCo environments we tested on.
@@ 4 Theoretical analysis
Consider the MDP $M=(\mathcal{S},\mathcal{A},P,c,\boldsymbol{\nu}_{0},\gamma)$ and two policies $\pi,\pi^{\prime}:\mathcal{S}\rightarrow\Delta_{\mathcal{A}}$. Then consider for any $\widehat{Q}\in\mathbb{R}^{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|}$ and $\widehat{V}^{\pi}(s)=\left\langle{\pi(\cdot|s)},{\widehat{Q}(s,\cdot)}\right\rangle$ and $Q^{\pi^{\prime}},V^{\pi^{\prime}}$ be respectively the state-action and state value function of the policy $\pi$ in MDP $M$. Then, it holds that $(1-\gamma)\left\langle{\boldsymbol{\nu}_{0}},{\widehat{V}^{\pi}-V^{\pi^{\prime}}}\right\rangle$ equals
@@
Consider an MDP where $\max_{s,a\in\mathcal{S}\times\mathcal{A}}\mathrm{supp}(P(\cdot|s,a))=2$, then for $L\geq 36\log\left({\frac{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|K}{\delta}}\right)$ it holds that with probability at least $1-\delta$
@@
Optimal choice of the number of critics network $L$
It is important to notice that Corollary 4.11 and Lemma 4.13 creates a tradeoff for what concerns the optimal choice of the number of critics. In particular, from Corollary 4.11, $L$ should be chosen large enough to ensure that optimism holds with high enough probability. On the other hand, one can notice that Lemma 4.13 upper bounds the expected on policy temporal difference error as $\mathcal{O}(L)$ therefore a smaller number of critics ensures a tighter bound. All in all, the best choice is the smallest $L$ that ensures optimism with probability at least $1-\delta$, that is $L=36\log\left({\frac{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|K}{\delta}}\right)$. The tradeoff with respect to the number of critics is also observed in a practical ablation study (see Figures 5 and 4) .
@@
To fix this issue, we consider multiple critics and we used as an optimistic estimate the mean minus the standard deviation of the ensemble as explained in Algorithm 5. In addition, the standard deviation needs to be truncated at a threshold, as was done in the tabular analysis, to avoid the value function estimators growing out of the attainable range. For any state $s$, the estimated value functions are truncated in the interval $\left[{0,(1-\gamma)^{-1}}\right]$.
@@ 6 Experiments
We highlight that the SOAR algorithmic idea can be used also for other imitation learning algorithms based on SAC such as AdRIL [90] and SMILING [108].
@@
6.1 Experiment on a hard exploration task
An anonymous reviewer pointed out that the MuJoCo benchmark is not the hardest for what concern exploration. This is a very valid suggestion that we address here. This will also allow to understand better the role of the number of critics $L$. Therefore, to highlight even more the importance of exploration especially in imitation learning from states only we run SOAR-IL in the worst case construction used in the lower bound for the number of environment interaction in [52, Theorem 19]. This is a simple two states MDP ( a low reward state and a high reward state ) with 20 actions per state. From the high reward state all actions are identical. From the low reward state, all actions are identical but the one chosen by the deterministic expert which has just a slightly higher probability to lead to the high reward state from the low reward state. Even observing the expert state occupancy measure perfectly, it is difficult for the learner to find out which is the action which the expert took. That is because all actions are almost identical but one.
Figure 4: Ablation for $L$ on hard exploration task. State only imitation experiment in a hard exploration environment (used in the lower bound from [52, Theorem 19]) . Results averaged over $5$ seeds, for a dataset of $100$ states sampled from the expert occupancy measure.
We can see that with only 1 network, the mean of the learner does not reach the expert performance and the variance is very high meaning that some seeds are successful and others fail. This is in perfect agreement with Corollary 4.11 which predicts that for low values of the number of critics $L$, the optimistic properties of the critic estimators can not be guaranteed with high probability. For $L=2$, the environment is solved successfully albeit with a higher variance than the case $L=3$. Increasing $L$ further leads to worst results in terms of MDP samples needed to solve the task. This is because according to Lemma 4.13 the upper bound on the expected on policy temporal difference error scales with $L$ so an excessively large $L$ should be avoided. We used $\alpha=0.5$, $\eta=4$, and we scaled the standard deviation bonus by $0.001$.
@@ Appendix A Related Works
The first theoretical guarantees obtained for the imitation learning problem dates back to the work of [2] and [93] which notably used the idea of no-regret learning. However, their work requires either knowledge of the environment transitions or they require a suboptimal in the precision parameter $\varepsilon$ amount of expert trajectories to estimate those. Our theoretical guarantees are in the same setting of previous works like [85] and [110] which do not require knowledge of the environment transitions a priori but assumes online trajectory access to the environment. The main difference that their work focuses on the easier finite horizon setting. Additionally, their exploration techniques only applies to the tabular setting. Indeed, in the MuJoCo experiments in [85], the authors do not attempt to implement the exploration mechanism required for their theoretical guarantees. In a similar way, also [110] can not be implemented beyond the tabular setting because it relies on a reward free procedure requiring bonuses proportional to the number of visits to each state action pair. [79] suggest an algorithm which does not require exploration but it can not improve upon behavioural cloning in terms of expert trajectories. [74, 76, 20] analyze instead offline imitation learning (behavioural cloning) where no additional interaction with the environment is allowed. This setting is more general but it comes at the cost of additional assumptions such as policy realizability or worst depedence on the horizon on the required number of trajectories. [20] presents an analysis for general policy classes but they require a maximum likelihood oracle which can not be implemented exactly when using neural network function approximation.
@@
As mentioned only few works investigated exploration techniques in Deep IL. Apart from the previously mentioned works, [114] adopts a model based approach and used exploration bonuses based on prediction error of the next observed state (a.k.a. curiosity driven exploration [68, 7]). Finally we notice that ensembles have been used in IL theory IL also for goals different to exploration. In particular, [91] partitioned the expert dataset in two subdataset and show that these technique allows for improved expert sample complexity bounds when the expert is deterministic.
@@ B.1 Upper bounding the policy regret
where the last inequality holds for $N^{k}_{\ell}\geq 2$. Otherwise, for $P(s^{\prime}|s,a)\leq 1-\frac{1}{N^{k}_{\ell}(s,a)}$ we can apply [9, Lemma 2] (adapted from [107, Corollary 1] ) to obtain
@@ B.2 Policy Regret Decomposition
$\sum^{K}_{k=1}\sum_{s,a}\left[{d^{\pi^{k}}(s,a)-d^{\pi^{\star}}(s,a)}\right]\cdot\left[{\delta^{k}(s,a)}\right]$ $\leq\gamma\sum^{K}_{k=1}\mathbb{E}_{s,a\sim d^{\pi^{k}}}\left[{\sqrt{\frac{\left|{\mathcal{S}}\right|}{(N^{k}(s,a)/L+1)(1-\gamma)^{2}}\log\left({\frac{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|LK^{2}(K+1)}{(1-\gamma)\delta}}\right)}}\right]$
@@
$\leq\sqrt{K\sum^{K}_{k=1}\mathbb{E}_{s,a\sim d^{\pi^{k}}}\left[{\frac{\left|{\mathcal{S}}\right|}{(N^{k}(s,a)/L+1)(1-\gamma)^{2}}\log\left({\frac{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|LK^{2}(K+1)}{(1-\gamma)\delta}}\right)}\right]}$
@@
$\sum^{K}_{k=1}\sum_{s,a}$ $\left[{d^{\pi^{k}}(s,a)-d^{\pi^{\star}}(s,a)}\right]\cdot\left[{\delta^{k}(s,a)}\right]$
$\leq\frac{1}{1-\gamma}\sqrt{K\left|{\mathcal{S}}\right|\log\left({\frac{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|LK^{2}(K+1)}{(1-\gamma)\delta}}\right)(2L\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|\log\left({KL_{\max}}\right)+4\log\left({2K/\delta}\right))}$
@@
$\leq\left({\sqrt{\frac{\left|{\mathcal{S}}\right|}{(N^{k}(s,a)/L+1)(1-\gamma)^{2}}\log\left({\frac{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|LK^{2}(K+1)}{(1-\gamma)\delta}}\right)}+\frac{2}{K}+\frac{2}{(N^{k}(s,a)/L+1)(1-\gamma)}}\right)^{2}$
$\leq\frac{3\left|{\mathcal{S}}\right|}{(N^{k}(s,a)/L+1)(1-\gamma)^{2}}\log\left({\frac{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|LK^{2}(K+1)}{(1-\gamma)\delta}}\right)+\frac{12}{K^{2}}+\frac{12}{(N^{k}(s,a)/L+1)^{2}(1-\gamma)^{2}}$
@@
$PV^{k}(s,a)-\widehat{P}^{k}_{\ell}V^{k}(s,a)$ $\leq\sqrt{\frac{\left|{\mathcal{S}}\right|}{(N^{k}(s,a)/L+1)(1-\gamma)^{2}}\log\left({\frac{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|LK^{2}(K+1)}{(1-\gamma)\delta}}\right)}+\frac{2}{K}$
$\phantom{=}+\frac{2}{(N^{k}(s,a)/L+1)(1-\gamma)}~{}~{}~{}~{}~{}\forall\ell,k\in[L]\times[K]$
In particular, the above statement implies that for all $k\in[K]$
$PV^{k}(s,a)-\min_{\ell\in[L]}\widehat{P}^{k}_{\ell}V^{k}(s,a)$ $\leq\sqrt{\frac{\left|{\mathcal{S}}\right|}{(N^{k}(s,a)/L+1)(1-\gamma)^{2}}\log\left({\frac{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|LK^{2}(K+1)}{(1-\gamma)\delta}}\right)}+\frac{2}{K}$
@@
Then, notice that denoting $s^{\prime}_{n}(s,a)$ the state sample after $s,a$ the $n^{th}$ time the state action pair was visited we have that
$\sum_{\bar{s},\bar{a},s^{\prime}\in\mathcal{R}^{k}_{\ell}}\left({P\tilde{V}(s,a)-\tilde{V}(s^{\prime})}\right)\mathds{1}_{\left\{{s,a=\bar{s},\bar{a}}\right\}}=\sum^{N^{k}_{\ell}(s,a)}_{n=1}\left({P\tilde{V}(s,a)-\tilde{V}(s^{\prime}_{n}(s,a))}\right)$
Applying directly the Azuma Hoeffding inequality is not possible because the number of elements in the sum, i.e. the number of visits $N^{k}_{\ell}(s,a)$ is not a random variable independent on the random variables $\left\{{s^{\prime}_{n}(s,a)}\right\}^{N^{k}_{\ell}(s,a)}_{n=1}$ (see [44, Exercise 7.1] ).
Therefore, we first apply the Azuma Hoeffding inequality for a specific $k$ and for a specific value of the visits $N^{k}_{\ell}(s,a)$. That is, it holds that with probability $1-\delta$
$\sum_{\bar{s},\bar{a},s^{\prime}\in\mathcal{R}^{k}_{\ell}}\left({P\tilde{V}(s,a)-\tilde{V}(s^{\prime})}\right)\mathds{1}_{\left\{{s,a=\bar{s},\bar{a}}\right\}}$ $\leq\sqrt{\frac{N^{k}_{\ell}(s,a)\log(1/\delta)}{2(1-\gamma)^{2}}}$
Therefore via a union bound for $k\in[K]$ and $N^{k}_{\ell}(s,a)\in\left\{{0,1,\dots,K}\right\}$ we have that with probability $1-\delta$ it holds that for all $k\in[K]$
$\sum_{\bar{s},\bar{a},s^{\prime}\in\mathcal{R}^{k}_{\ell}}\left({P\tilde{V}(s,a)-\tilde{V}(s^{\prime})}\right)\mathds{1}_{\left\{{s,a=\bar{s},\bar{a}}\right\}}$ $\leq\sqrt{\frac{N^{k}_{\ell}(s,a)\log(K(K+1)/\delta)}{2(1-\gamma)^{2}}}.$
Therefore, we can conclude that with probability at least $1-\delta$ for all $k\in[K]$
$P\tilde{V}(s,a)-\frac{1}{N^{k}_{\ell}(s,a)}\sum_{\bar{s},\bar{a},s^{\prime}\in\mathcal{R}^{k}_{\ell}}\tilde{V}(s^{\prime})\mathds{1}_{\left\{{s,a=\bar{s},\bar{a}}\right\}}\leq\sqrt{\frac{\log(K(K+1)/\delta)}{2N^{k}_{\ell}(s,a)(1-\gamma)^{2}}}.$
Now, by a another union bound over $\mathcal{C}_{\epsilon_{\mathrm{cov}}}(\mathcal{V})$ , $[K]$, $[L]$ and $\mathcal{S}\times\mathcal{A}$ and denoting
@@
$\mathbb{P}\left[{P\tilde{V}(s,a)-\bar{P}^{k}_{\ell}\tilde{V}(s,a)\leq\sqrt{\frac{\log(K(K+1)\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|\left|{\mathcal{C}_{\epsilon_{\mathrm{cov}}}(\mathcal{V})}\right|L/\delta)}{N^{k}_{\ell}(s,a)(1-\gamma)^{2}}}~{}~{}\forall s,a,\ell,k\in\mathcal{S}\times\mathcal{A}\times[L]\times[K],~{}~{}\tilde{V}\in\mathcal{V}}\right]\geq 1-\delta$
@@
$\leq\sqrt{\frac{\log(K(K+1)\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|\left|{\mathcal{C}_{\epsilon_{\mathrm{cov}}}(\mathcal{V})}\right|L/\delta)}{N^{k}_{\ell}(s,a)(1-\gamma)^{2}}}+2\epsilon_{\mathrm{cov}}$
$\leq\sqrt{\frac{\left|{\mathcal{S}}\right|}{N^{k}_{\ell}(s,a)(1-\gamma)^{2}}\log\left({\frac{K(K+1)\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|L}{(1-\gamma)\epsilon_{\mathrm{cov}}\delta}}\right)}+2\epsilon_{\mathrm{cov}}$
@@
$PV^{k}(s,a)-\bar{P}^{k}_{\ell}V^{k}(s,a)\leq\sqrt{\frac{\left|{\mathcal{S}}\right|}{N^{k}_{\ell}(s,a)(1-\gamma)^{2}}\log\left({\frac{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|LK^{2}(K+1)}{(1-\gamma)\delta}}\right)}+\frac{2}{K}$
@@
$\geq\frac{N^{k}_{\ell}(s,a)}{N^{k}_{\ell}(s,a)+2}\left[{PV^{k}(s,a)-\sqrt{\frac{\left|{\mathcal{S}}\right|}{N^{k}_{\ell}(s,a)(1-\gamma)^{2}}\log\left({\frac{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|LK^{2}(K+1)}{(1-\gamma)\delta}}\right)}-\frac{2}{K}}\right]$
$=PV^{k}(s,a)-\frac{2}{N^{k}_{\ell}(s,a)+2}PV^{k}(s,a)-\sqrt{\frac{\left|{\mathcal{S}}\right|}{(N^{k}_{\ell}(s,a)+2)(1-\gamma)^{2}}\log\left({\frac{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|LK^{2}(K+1)}{(1-\gamma)\delta}}\right)}-\frac{2}{K}$
$\geq PV^{k}(s,a)-\frac{2}{(N^{k}_{\ell}(s,a)+2)(1-\gamma)}-\sqrt{\frac{\left|{\mathcal{S}}\right|}{(N^{k}_{\ell}(s,a)+2)(1-\gamma)^{2}}\log\left({\frac{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|LK^{2}(K+1)}{(1-\gamma)\delta}}\right)}-\frac{2}{K}$
Finally, rearranging and using that $N^{k}_{\ell}(s,a)=\left\lfloor\frac{N^{k}(s,a)}{L}\right\rfloor\geq\frac{N^{k}(s,a)}{L}-1,$ we obtain that with probability $1-\delta$, it holds that for all $\ell\in[L]$, $k\in[K]$, $s,a\in\mathcal{S}\times\mathcal{A}$
$PV^{k}(s,a)-\widehat{P}^{k}_{\ell}V^{k}(s,a)$ $\leq\sqrt{\frac{\left|{\mathcal{S}}\right|}{(N^{k}(s,a)/L+1)(1-\gamma)^{2}}\log\left({\frac{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|LK^{2}(K+1)}{(1-\gamma)\delta}}\right)}+\frac{2}{K}$